A 50.0 mL sample of 0.436 M NH_4NO_3 is diluted wi
Answer #1

from dilution formula

M1V1 = M2V2

50*0.436 = 250*M2

M2 = 0.0872 M

M1 = molarity of standard solution = 0.436 M

V1 = volume of standard solution =50 ml

V2 = volume of resulting solution= 250 ml

M2 = molarity of resulting solution = 0.0872 M

    = 8.72*10^-2 M